### v1

# Inflection Point Inflation in Supergravity

Abstract
In this paper, we study the inflection point inflation generated by a polynomial superpotential and a canonical Kähler potential under the supergravity framework, where only one chiral superfield is needed. We find that the special form of the scalar potential limits the inflationary Hubble parameter to values $\lesssim 10^{10}\,\textrm{GeV}$ and the inflaton mass to $\lesssim 10^{11}\,\textrm{GeV}$. We obtain analytic results for small field cases and present numerical results for large field ones. We find the tensor-to-scalar ratio $r<10^{-8}$ is always suppressed in these models, while the running of spectral index $\alpha\approx\mathcal{O}(-10^{-3})$ may be testable in next-generation CMB experiments. We also discuss the possible effects of SUSY breaking Polonyi term presented in the superpotential where we find a general upper bound for the SUSY breaking scale for a given value of the Hubble parameter.
1 Introduction
Recent observations of the cosmic microwave background (CMB) by the Planck and BICEP/Keck experiments strongly favor an exponentially expanding period of the early universe Planck:2018jri ; BICEP:2021xfz , which is called inflation Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi . The simplest implementation, slow roll inflation, needs a single scalar field “inflaton” $\phi$ rolls over a flat region in the potential. The flatness of the potential can be parameterized by the following slow roll parameters [^1] We set the reduced Planck Mass $M_{\textrm{pl}}=\sqrt{\frac{1}{8\pi G}}\approx 2.4\times 10^{18}\,\textrm{GeV}=1$ throughout this paper. Lyth:2009zz :
$\epsilon=\frac{1}{2}\left(\frac{V^{\prime}}{V}\right)^{2}\,,\quad\eta=\frac{V^{\prime\prime}}{V}\,,\quad\xi^{2}=\frac{V^{\prime}V^{\prime\prime\prime}}{V^{2}}\,.$ (1)
where a prime () denotes the derivative with respect to $\phi$. They are related with the CMB observations Planck:2018jri ; BICEP:2021xfz :
$\begin{split}P_{\zeta}&=\frac{V}{24\pi^{2}\epsilon}=(2.1\pm 0.1)\times 10^{-9}\,;\\ n_{s}&=1-6\epsilon+2\eta=0.9659\pm 0.0040\,;\\ r_{0.05}&=16\epsilon<0.036\,;\\ \alpha&=16\epsilon\eta-24\epsilon^{2}-2\xi^{2}=-0.0041\pm 0.0067\,.\end{split}$ (2)
Here $P_{\zeta}$ is the power spectrum of Gaussian curvature perturbations, $n_{s}$ is the spectral index, $r$ is the tensor-to-scalar ratio and $\alpha$ is is the running of the spectral index.
Power law potentials $V\propto\phi^{n}$ offer one of the most elegant and simplest realizations of inflation. However, such potentials with $n>1$ have been ruled out by the CMB observations, due to the large tensor-to-scalar ratio predicted by these models. Given this fact, a broad program has been pursed to investigate alternative models for inflation. Among them are the so-called inflection point inflation models, where different $\phi^{n}$ terms collectively provide an inflection point in the potential and accommodate inflation around it. Inflection point models have attracted much attention in recent years  Enqvist:2010vd ; Mazumdar:2011ih ; Hotchkiss:2011am ; Chang:2013cba ; Gao:2015yha ; Okada:2016ssd ; Gao:2016xfv ; Choi:2016eif ; Dimopoulos:2017xox ; Okada:2017cvy ; Pallis:2023aom . They have been well studied both during and after inflation Drees:2021wgd ; Drees:2022aea ; Drees:2024hok ; Bernal:2024ykj ; Drees:2025iue .
Loop corrections can easily spoil the flatness of the inflaton potential, while supersymmetry (SUSY) provides a convenient way to protect it. For this reason, different inflation models have often been considered in a supergravity (SUGRA) framework Nanopoulos:1982bv ; Ovrut:1983my ; Holman:1984yj ; Adams:1992bn ; Copeland:1994vg ; Kumekawa:1994gx ; Izawa:1996dv ; Linde:1997sj ; Binetruy:2004hh ; Allahverdi:2006we ; Ben-Dayan:2008fhy ; Ellis:2013xoa ; Ellis:2013nxa ; Ferrara:2013rsa ; Kallosh:2014rga ; Ketov:2014qha ; Carrasco:2015pla ; Romao:2017uwa . This paper will follow the same spirit to consider a supersymmetric inflection point inflaton model. We want to present a systematic study allowing us to scan over the possible parameter space. We will focus on the similarities and differences between the previously studied non-supersymmetric, renormalizable models and SUGRA models.
SUSY must be broken since no superpartner of a Standard Model particle has as yet been detected. Once SUSY breaking terms are included, they may modify the form of the inflaton potential, thereby breaking the slow-roll conditions. Thus, the SUSY breaking scale must be bounded from above for successful inflation. To investigate this effect, we will focus on the case where SUSY is broken by the Polonyi model Polonyi:1977pj . We find two limiting cases by comparing the relative strength of the inflation sector and the SUSY breaking sector. They lead to very different bounds on the SUSY breaking scale.
To be more precise, we consider a SUSY-preserving inflaton field $\phi$, which accommodates a near inflection point in the scalar potential at $\phi=\phi_{0}$. We find that the Hubble parameter and the inflaton mass both increase with $\phi_{0}$ when $\phi_{0}<1$. They reach a maximum around $\phi_{0}\sim 1$, then start to decrease because of the exponential factor in the potential. Thus, in such a model, the Hubble parameter can not exceed $\mathcal{O}(10^{10})\textrm{GeV}$. The tensor-to-scalar ratio $r$ is much smaller than the current upper bound. The model also predicts a near constant running $\alpha\sim-0.003$, a unique feature that the next generation of observations might be able to test.
The remainder of this paper is organized as follows. In sec. 2 we recap basic results obtained in the non-supersymmetric, renormalizable version of the model and argue why it is interesting to consider this scenario in the SUGRA framework. In sec. 3 we present our model’s analytic and numerical results. We further discuss the SUSY breaking effects in our model. In sec. 4 we summarize the work and draw some conclusions.
2 Renormalizable Inflection Point Model and Beyond
2.1 Potential setup and CMB observables
We start from the most general renormalizable potential of a single real scalar inflaton field $\phi$:
.$\,.V(\phi)=b\phi^{2}+c\phi^{3}+d\phi^{4}\,,$ (3)
where we have removed the constant and linear terms, so that the minimum of the potential is defined to be at $\phi=0$ with vanishing vacuum energy [^2] A tiny cosmological constant can be added.. Requiring an inflection point, $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$, then leads to:
$b=\frac{9c^{2}}{32d}\,,\quad\phi_{0}=-\frac{3c}{8d}\,.$ (4)
CMB observations indicate the potential is not exactly flat, but rather concave. One way to realize this is to introduce a small deviation from the inflection point conditions in the cubic term. For our purpose it is more convenient to write the coefficients in terms of the inflection point position $\phi_{0}$. The modified potential then reads:
$V(\phi)=d\left(\phi^{4}-\frac{8}{3}\phi_{0}(1-\beta)\phi^{3}+2\phi_{0}^{2}\phi^{2}\right)\,.$ (5)
There are three free parameters in the potential, $d,\phi_{0}$ and $\beta$. In eq.(5), $d$ determines the overall normalization of the potential, which can be matched to the power spectrum of curvature perturbation $P_{\zeta}$ once the other parameters are fixed. The other two parameters govern the shape the potential and hence determine the CMB observables, such as the number of e-folds when the CMB pivot scale left the horizon, $N_{\textrm{cmb}}$, and the spectral index $n_{s}$. In a “small field” set–up, i.e. for sub-Planckian field values, $\phi_{0}<1$, fixing $N_{\textrm{cmb}}=65$ and eq.(2) require Drees:2021wgd :
$\begin{split}d&=6.61\times 10^{-16}\phi_{0}^{2}\,;\\ \beta&=9.73\times 10^{-7}\phi_{0}^{4}\,.\end{split}$ (6)
This leads to the following predictions:
$\begin{split}b&=1.3\times 10^{-15}\,\phi_{0}^{4}\,;\quad c=1.8\times 10^{-15}\,\phi_{0}^{3}\,;\\ \indent H_{\textrm{inf}}&=8.6\times 10^{-9}\,\phi_{0}^{3}\,;\quad m_{\phi}=5.2\times 10^{-8}\,\phi_{0}^{2}\,;\\ \indent r&=7.1\times 10^{-9}\,\phi_{0}^{6}\,;\quad\alpha=-1.4\times 10^{-3}\,.\end{split}$ (7)
The same ansatz for the potential can also describe “large field” scenarios, where $\phi_{0}>1$. However, in this case no analytical treatment is known. Numerical studies showed that this model can cover the whole allowed parameter space in the $n_{s}-r$ plane and may even lead to double eternal inflation Drees:2022aea .
2.2 Possible Realization in SUGRA and Associated Problem
In SUGRA the scalar potential is generated by the Kähler potential $K$ and the holomorphic Superpotential $W$ [^3] For a detailed introduction, please consult Freedman:2012zz .. Both of them are supposed to be functions of the complex field $\Phi$, which is the scalar component of a chiral superfield. The Kähler potential also determines the kinetic term of the scalar field:
$\mathcal{L}_{\textrm{kin}}=-\frac{\partial^{2}K}{\partial\Phi\partial\Phi^{*}}\partial_{\mu}\Phi\partial^{\mu}\Phi^{*}\,.$ (8)
A canonically normalized field thus requires $K=-\Phi\Phi^{*}$ and we will use this Kähler potential throughout the paper. We can further define the Kähler covariant derivative as:
$D_{\Phi}W=\frac{\partial W}{\partial\Phi}+\frac{\partial K}{\partial\Phi}W\,.$ (9)
The $F-$term contribution to the scalar potential then reads:
[…]
Since we assume the inflaton to be a gauge singlet, there is no $D-$term contribution.
For simplicity we choose the superpotential to be a polynomial function of the complex field $\Phi$:
$W(\Phi)=B\Phi^{2}+C\Phi^{3}+D\Phi^{4}\,,$ (11)
where we neglect the constant and linear terms as in the renormalizable case. The ansatz (11) ensures that $V(0)=W(0)=0$, i.e. the potential has a supersymmetric stationary point at $\Phi=0$. The three coefficients $B,C,D$ are chosen to be real. The complex $\Phi$ can in general be written as $\Phi=(\phi+i\chi)/\sqrt{2}$, where $\phi$ and $\chi$ are real fields. We want to identify $\phi$ with the inflaton field. By choosing the coefficients in eq.(11) to be real we make sure that the potential $V(\phi,\chi)$ depends only on even powers of $\chi$. Moreover, we make sure that $\partial^{2}V(\phi,\chi)/\partial\chi^{2}>0$ for $\chi=0$ and $\phi\in[0,\phi_{0}]$. We therefore can assume that $\chi=0$ throughout, so that we can ignore this imaginary component when computing the inflationary dynamics. It is not hard to write down the scalar potential $V(\phi)\equiv V(\phi,\chi=0)$ in this setup:
$\begin{split}V(\phi)={\rm e}^{\frac{\phi^{2}}{2}}\Biggl{[}2B^{2}\phi^{2}+3\sqrt{2}BC\phi^{3}+\frac{1}{4}\phi^{4}\left(B^{2}+16BD+9C^{2}\right)+\frac{1}{2}\phi^{5}\left(\sqrt{2}BC+6\sqrt{2}CD\right)\\ \indent+\frac{1}{8}\phi^{6}\left(B^{2}+6BD+3C^{2}+16D^{2}\right)+\frac{1}{8}\phi^{7}\left(\sqrt{2}BC+4\sqrt{2}CD\right)\\ \indent+\frac{1}{16}\phi^{8}\left(2BD+C^{2}+5D^{2}\right)+\frac{CD\phi^{9}}{8\sqrt{2}}+\frac{D^{2}\phi^{10}}{32}\Biggr{]}\,,\end{split}$ (12)
In order to try to match to the renormalizable case, we expand $V$ up to the fourth order:
$V(\phi)\approx 2B^{2}\phi^{2}+3\sqrt{2}BC\phi^{3}+\frac{1}{4}(5B^{2}+9C^{2}+16BD)\phi^{4}\,.$ (13)
Matching eq.(13) to eq.(3) and using eqs.(6) and (7) implies:
$\begin{split}B&=2.571\times 10^{-8}\phi_{0}^{2}\,,\\ C&=1.615\times 10^{-8}\phi_{0}\times(1-\beta)\,,\\ D&=7.210\times 10^{-10}-8.034\times 10^{-9}\phi_{0}^{2}-5.7\times 10^{-8}(2\beta-\beta^{2})\,.\end{split}$ (14)
Hence for $\phi_{0}\ll 1$ the coefficients of the superpotential would have to scale as $B\propto\phi_{0}^{2},\ C\propto\phi_{0}^{1},\ D\propto\phi_{0}^{0}$. The problem emerges when we substitute these matching conditions back into the full potential eq.(12). Writing the latter as:
$V(\phi)=e^{\frac{1}{2}\phi^{2}}\sum_{n=2}^{10}a_{n}\phi^{n}\,,$ (15)
the coefficients $a_{n}$ have the following scaling behavior for $\phi_{0}\ll 1$:
$\begin{split}&a_{2}\propto\phi_{0}^{4};\quad a_{3}\propto\phi_{0}^{3};\quad a_{4}\propto\phi_{0}^{2};\quad a_{5}\propto\phi_{0}^{1};\quad a_{6}\propto\phi_{0}^{0};\\ \indent&a_{7}\propto\phi_{0}^{1};\quad a_{8}\propto\phi_{0}^{0};\quad a_{9}\propto\phi_{0}^{1};\quad a_{10}\propto\phi_{0}^{0}.\end{split}$ (16)
Thus, when we evaluate the value of the potential and its first and second derivatives at $\phi\sim\phi_{0}$, the first five terms (from $\phi^{2}$ to $\phi^{6}$) would contribute with comparable magnitude. Therefore the terms $\propto\phi^{5}$ and $\phi^{6}$ can easily spoil the flatness of the potential, i.e. the expansion only to order $\phi^{4}$ is not self–consistent. For this reason, it is necessary to consider an inflection point model in the full potential eq.(12) (or at least up to $\phi^{6}$) rather than trying to directly match the renormalizable, non-supersymmetric potential.
3 Inflection Point Model in SUGRA
3.1 Analytic Analysis of the Model
Requiring that the full potential eq.(12) has an inflection point at $\phi=\phi_{0}$, i.e. $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$, leads to the following solutions [^4] Since we are dealing with high-order polynomial equations there often are several solutions. However, we find that the others usually have $V<0$ at the minimum, leading to a very large and negative cosmological constant.:
$\begin{split}B=&D\frac{\phi_{0}^{2}\left(1152+480\phi_{0}^{2}+72\phi_{0}^{4}+12\phi_{0}^{6}+\phi_{0}^{8}\right)}{2\left(192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}\right)}\\ C=&-D\frac{\sqrt{2}\phi_{0}\left(384+192\phi_{0}^{2}+24\phi_{0}^{4}+8\phi_{0}^{6}+\phi_{0}^{8}\right)}{192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}}\end{split}$ (17)
For $\phi_{0}\ll 1$ we can approximate the above solutions by their leading order results:
$B\approx D\times 3\phi_{0}^{2}\,,\quad C\approx D\times(-2\sqrt{2}\phi_{0})\,.$
The full potential at $\phi\leq\phi_{0}$ can be further simplified if we only include terms up to $\phi_{0}^{6}$:
$V(\phi)\approx 2D^{2}\phi^{2}(\phi^{2}-3\phi\phi_{0}+3\phi_{0}^{2})^{2}\,.$ (18)
This expansion now is self-consistent, i.e. the higher order terms, starting at ${\cal O}(\phi^{7})$, are indeed suppressed. At the inflection point $\phi_{0}$, the potential reads:
$V(\phi_{0})\approx 2D^{2}\phi_{0}^{6}\,.$ (19)
The potential (18) has a minimum at $\phi=0$ and is positive semi-definite, i.e. $V(\phi)\geq 0\ \forall\phi$. As in the non-supersymmetric case, in the small field scenario $\phi_{0}\ll 1$ we can get semi-analytic results for inflationary observables by expanding the slow-roll parameter around the inflection point via the ansatz $\phi=\phi_{0}(1-\delta\phi)$, and adding a deviation from the strict inflection point condition via $B\to B+D\times\delta B$:
$\begin{split}\epsilon&=\frac{1}{2}\left(\frac{V^{\prime}}{V}\right)^{2}=\frac{1}{2}\left(\frac{2\delta B+6\phi_{0}^{2}\delta\phi^{2}}{\phi_{0}^{3}}\right)^{2}\,,\\ \eta&=\frac{V^{\prime\prime}}{V}=\frac{10.5\,\delta B-24\,\delta\phi}{2\phi_{0}^{2}}\,,\\ \xi^{2}&=\frac{V^{\prime}V^{\prime\prime\prime}}{V^{2}}=\frac{24\delta B+72\phi_{0}^{2}\delta\phi^{2}}{\phi_{0}^{6}}\,.\end{split}$ (20)
We only keep terms up to linear order in $\delta B$ and quadratic in $\delta\phi$; this will turn out to be sufficient.
We are now ready to discuss which $\delta\phi$ and $\delta B$ reproduce the CMB observations. In our model the duration of inflation is controlled by $\eta$ since $\epsilon\ll\eta$. The beginning (when the pivot scale $k_{*}=0.05\,\textrm{Mpc}^{-1}$ crossed out of the horizon) and end of observable inflation are given by:
$\eta_{\textrm{cmb}}=\frac{n_{s}-1}{2}\,,\quad\eta_{\textrm{end}}=-1\,,$ (21)
Solving the second eq.(20) for $\delta\phi$, we get:
$\delta\phi=-\frac{2\phi_{0}^{2}\eta-10.5\delta B}{24}\,.$ (22)
The number of e-folds $N_{\textrm{cmb}}$ is given by:
[…]
In the last step we have switched from the absolute deviation $\delta B$ to the relative one $\beta=D\times\delta B/B=\delta B/(3\phi_{0}^{2})$ in order to factor out $\phi_{0}$ in the denominator; except for the first line, eq.(23) also only holds for $\phi_{0}\ll 1$. Recall that the arc-tangent functions can be at most $\pi/2$. Numerically, requiring $N_{\textrm{cmb}}=50$ yields $\beta=2.7\times 10^{-5}\phi_{0}^{4}$ and $\delta B=8.2\times 10^{-5}\phi_{0}^{6}$. From the second eq.(2) and remembering $|\epsilon|\ll|\eta|$ we see that $|\eta|\geq 0.015$ when and after CMB scales crossed out of the horizon. Hence we can neglect the second term in eq.(22) for $\phi_{0}\ll 1$, and determine the inflation period by:
$\delta\phi=-\frac{\phi_{0}^{2}}{12}\eta\,.$ (24)
Inserting this into the last line of eq.(23) yields
$N_{\textrm{cmb}}=\frac{\phi_{0}^{2}}{6\sqrt{\beta}}\left(\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{end}}}{12\sqrt{\beta}}}\right)-\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{cmb}}}{12\sqrt{\beta}}}\right)\right)\,,$ (25)
which can be solved numerically. For example, setting $n_{s}=0.9659\,,N_{\textrm{cmb}}=45$ would result in $\delta B=6.1\times 10^{-5}\phi_{0}^{6}$.
Having fixed $\delta B$ and the initial $\delta\phi$, the overall scale of inflation, and hence $D$, is determined by the power of Gaussian curvature perturbations. To this end we first calculate $\epsilon$ at $\phi_{\textrm{cmb}}$:
$\epsilon_{\textrm{cmb}}=8.88\,\times 10^{-9}\phi_{0}^{6}\,,$ (26)
Using eqs.(2) and (19) we find that the normalization factor is independent of $\phi_{0}$:
$D^{-2}=\frac{2\phi_{0}^{6}}{P_{\zeta}24\pi^{2}\epsilon_{\textrm{cmb}}}=4.52\times 10^{14}\,.$ (27)
It is then straightforward to evaluate the value of the Hubble parameter during inflation and the physical mass of the inflaton after inflation:
$H_{\textrm{inf}}=\sqrt{\frac{V(\phi_{0})}{3}}=3.84\times 10^{-8}\phi_{0}^{3}\,,\quad m_{\phi}=\sqrt{4B^{2}}=2.81\times 10^{-7}\phi_{0}^{2}\,.$ (28)
The running of the spectral index $\alpha$ can also be determined:
$\alpha=16\epsilon\eta-24\epsilon^{2}-2\xi^{2}\approx-2\xi^{2}=-2\frac{24\delta B+72\phi_{0}^{2}\delta\phi^{2}}{\phi_{0}^{6}}\approx-0.0030\,.$ (29)
This running of the spectral index is a feature of our model, since it does not depend on $\phi_{0}$.
However, our numerical results do depend on $N_{\textrm{cmb}}$. Taking as second example the rather large value $N_{\textrm{cmb}}=65$ while keeping $n_{s}=0.9659$, we get:
$\begin{split}\delta B=2.60\times 10^{-5}\phi_{0}^{6}\,&,\quad\epsilon_{\textrm{cmb}}=1.59\times 10^{-9}\phi_{0}^{6}\,,\\ D^{-2}=2.53\times 10^{15}\,&,\quad H_{\textrm{inf}}=1.62\times 10^{-8}\phi_{0}^{3}\,,\\ m_{\phi}=1.19\times 10^{-7}\phi_{0}^{2}\,&,\quad\alpha=-0.0015\,.\end{split}$ (30)
The tensor-to-scalar ratio $r=16\epsilon$ is always too small to reach the sensitivity of any currently conceivable observation. However, the S4CMB experiment, together with small-scale structure information (e.g. on the Lyman-$\alpha$ forest) could achieve $10^{-3}$ sensitivity for $\alpha$, which would test our model Munoz:2016owz .
By comparing with the predictions of the renormalizable model given in eqs.(7), we see that of $H_{\textrm{inf}}$ and $m_{\phi}$ scale with $\phi_{0}$ in the same way in both cases, with roughly a factor $2$ difference in coefficients. Even though the SUGRA potential is more complicated and scales as $\phi^{6}$ in the simplest limit, they thus make very similar predictions for the inflationary observables. In particular, the overall coefficient $D$ of the sixth order SUGRA potential is independent of $\phi_{0}$, while in the non-SUSY case the coefficient of the quartic potential $d\propto\phi_{0}^{2}$; hence in both cases $V(\phi\simeq\phi_{0})\propto\phi_{0}^{6}$.
We also note that the curvature of the potential is negative for an extended range of $\phi$ below $\phi_{0}$. In the non-SUSY case, this holds for $\phi/\phi_{0}\in[1/3,1]$, independently of the value of $\phi_{0}$; in the SUGRA case with $\phi_{0}\ll 1$, this region extends to $\phi/\phi_{0}\in[1/4,1]$. The minimum of the curvature occurs at $\phi\simeq 0.54\phi_{0}$, closer to the origin than in the non-SUSY case, with a value just below $-0.23~{}m_{\phi}^{2}$, smaller in magnitude than in the non-SUSY case. The latter reduces the tachyonic instability while the former increases it. Hence, we expect the non-perturbative effects after inflation studied in Drees:2025iue for the non-SUSY case would be similar in the SUGRA version of the model.
When $\phi_{0}$ is larger than unity, it is hard to make a comprehensive analytic analysis, but we can still understand the model qualitatively. To this end we first formally rewrite the potential as:
$V(\phi)=e^{\frac{1}{2}\phi^{2}}P(\phi)\,.$ (31)
The slow roll parameter then scales like $\eta\propto\phi^{2}\times f(\delta\phi)$ with $f(0)\approx 0$. The duration of inflation is still controlled by $\delta\eta=\eta_{\textrm{cmb}}-\eta_{\textrm{end}}\approx 1$, the larger $\phi_{0}$, the smaller $\delta\phi$ should be. The resulting decrease in the integration range in eq.(23) has to be compensated by increasing the integrand, by a reduction of $\epsilon$. Since the ratio between the potential $V(\phi_{0})$ and $\epsilon$ is fixed by the power of curvature perturbations, an increase of $\phi_{0}$ would eventually lead to a decrease of the potential and hence of the Hubble parameter. Moreover, the potential near $\phi_{0}$ is increased by the exponential factor, which becomes unity at the origin. Hence, the mass of inflaton should be suppressed by $e^{-\frac{1}{2}\phi_{0}^{2}}$.
As a brief summary, by fixing $n_{s}$ and $N_{\textrm{cmb}}$, we find the inflection point model would always give a tiny tensor-to-scalar ratio $r$ and a constant running of spectral index $\alpha$. The Hubble scale of inflation first increases with increasing inflection point position $\phi_{0}$, and then decreases once $\phi_{0}$ exceeds $1$. The inflaton mass follows the same pattern but drops much faster in the second phase. In the next section, we confirm these expectations by showing some numerical results.
3.2 Numerical Results of the Model
In this section, we present our numerical results. We introduce three steps to scan the allowed parameter space of our model, following the same spirit as our analytic treatment:
  • We choose $\phi_{0}$ as a free parameter and solve the inflection point equations $V^{\prime}=V^{\prime\prime}=0$ to find corresponding values of $B/D$ and $C/D$. We pick the solution that will generate a positive semi-definite potential, see eqs.(17).
  • We slightly deform the potential by $B\to B+\delta B$. CMB scales start to leave the horizon at $\phi_{\textrm{cmb}}$, which is determined by $1+2\eta=n_{s}$ since still $\epsilon\ll|\eta|$ in all cases. Inflation ends at $\phi_{\textrm{end}}$, which is determined by $\eta=-1$. We find that both the start point and the end point mildly depend on $\delta B$. The correct $\delta B$ is given by fixing $N_{\textrm{cmb}}$, for which we consider the range from $45$ to $65$.
  • Having fixed $\delta B$, we can recalculate the slow roll parameters at the pivot scale $\eta_{\rm cmb}$ and $\epsilon_{\rm cmb}$. We then determine the correct normalization $D$ of the potential by requiring $P_{\zeta}=2.1\times 10^{-9}$. This allows us to compute the Hubble value and the inflaton mass.
We begin by showing four inflaton potentials with different choices of $\phi_{0}$ in Fig. 1. For comparison, these potentials are rescaled by their values at the inflection point. When $\phi_{0}<1$, the shape of the potential becomes independent of $\phi_{0}$ for $\phi\leq\phi_{0}$. Increasing $\phi_{0}$ beyond $1$ shortens the flat plateau; it also makes it even flatter, which is difficult to see in this figure.
[…]
The corresponding values of the Hubble parameter during inflation can be found in the top left frame of Fig. 2, where we use blue and orange lines to represent different values of $N_{\textrm{cmb}}$. As expected the Hubble scale first increases with increasing $\phi_{0}$, then drops once $\phi_{0}>1$. There is no lower bound on $H_{\rm inf}$ from the pure model perspective. However, the maximum value is determined by the special shape of the potential and can never exceed $10^{11}\ \textrm{GeV}$. It obeys a power law when $\phi_{0}$ is small, which agrees with our analytic estimation.
[…]
The running of spectral index $\alpha$ is shown in the bottom right frame. It remains independent of $\phi_{0}$ even for $\phi_{0}\geq 1$. In contrast to the non-SUSY version of the model, $\alpha$ strongly depends on $N_{\textrm{cmb}}$. When $N_{\textrm{cmb}}=45$ $\alpha\approx-0.0032$. Increasing $N_{\textrm{cmb}}$ reduces the absolute value of $\alpha$, reaching $\alpha\approx-0.0013$ for $N_{\textrm{cmb}}=65$.
Finally, the bottom left frame of Fig. 2 shows the relationship between the tensor-to-scalar ratio $r$ and $\phi_{0}$. It follows the same pattern as the Hubble scale. However, since we find $r<10^{-7}$, a positive detection of tensor modes by current or near-future experiments would exclude our model.
(a) Hubble scale $H_{\inf}$ during inflation.
(b) Inflaton mass $m_{\phi}$.
(c) Tensor to scale ratio $r$.
(d) Running of spectral index $\alpha$.
Figure 2: The dependence of the Hubble parameter $H_{\inf}$ during inflation, the inflaton mass $m_{\phi}$, tensor-to-scalar ratio $r$, and the running of spectral index $\alpha$ on the position $\phi_{0}$ of the inflection point. Different lines represent different choices of the number of e-folds: $N_{\textrm{cmb}}=65$ (blue) and $N_{\textrm{cmb}}=45$ (orange). We fixed $n_{s}=0.9659$ and $P_{\zeta}=2.1\times 10^{-9}$.
3.3 SUSY Breaking by a Polonyi Field
Clearly SUSY must be broken in any realistic model. In this subsection we investigate if the existence of a SUSY breaking sector would change the inflation potential significantly. For simplicity we consider the classical Polonyi ansatz Polonyi:1977pj , where a single chiral superfield $Z$ with a linear superpotential is introduced to break SUSY:
$\begin{split}W&=B\Phi^{2}+C\Phi^{3}+D\Phi^{4}+\mu M_{\textrm{pl}}(Z+\beta_{\rm P})\\ &=W_{\textrm{I}}+W_{\textrm{P}}\end{split}$ (32)
where $\mu$ essentially sets the SUSY breaking scale; we explicitly include a factor $M_{\textrm{pl}}$ here to ensure the correct dimension of $\mu$. Both $\Phi$ and $Z$ are complex fields. As before, we want the real part of $\Phi$ to be the inflaton field $\phi$. $Z$ is the Polonyi field whose vacuum expectation value $\langle Z\rangle$ is the only source of SUSY breaking after inflation, when $\langle\phi\rangle=0$. The SUSY breaking with vanishing vacuum energy requires $\beta_{\rm P}=2-\sqrt{3}$, which gives the gravitino mass $m_{3/2}=\mu e^{2-\sqrt{3}}$ when $Z$ stays at the SUSY breaking minimum at $Z=\langle Z\rangle=(\sqrt{3}-1)M_{\textrm{pl}}$.
Let’s first consider the case where the Polonyi sector gives a small perturbation to the inflation potential. For $\phi_{0}<1$ our previous results suggest:
$\begin{split}\tilde{B}&=\frac{B}{D}\approx 3\phi_{0}^{2}\,,\\ \tilde{C}&=\frac{C}{D}\approx-2\sqrt{2}\phi_{0}\,,\\ D&\approx 4.7\times 10^{-8}\,.\end{split}$ (33)
We require that the Polonyi field does not change the inflation potential significantly, which means $|W_{\textrm{I}}|\gg|W_{\textrm{P}}|$. The existence of the SUSY breaking term will not alter the slow roll parameters significantly if
$\epsilon_{\mu}\ll\epsilon_{\rm cmb}\quad{\rm and}\quad\eta_{\mu}\ll\eta_{\rm cmb}\,,$ (34)
where the subscript $\mu$ means the additional contribution to the slow roll parameter due to the SUSY breaking term.
We assume that during inflation the Polonyi field stays at the origin, $Z=0$, which will be verified later. Under this assumption, the additional contribution to the inflaton potential reads:
$V_{\mu}(\phi)=\frac{{\rm e}^{\frac{1}{2}\phi^{2}}}{4}\mu\left((-4\beta_{\rm P}B\phi^{2}+2\beta_{\rm P}(B+D)\phi^{4}+\sqrt{2}\beta_{\rm P}C\phi^{5}+\beta_{\rm P}D\phi^{6})+\mu(4-12\beta_{\rm P}^{2}+2\beta_{\rm P}^{2}\phi^{2})\right)\,.$ (35)
After substituting $\beta_{\rm P}=2-\sqrt{3}$, the additional contributions to slow roll parameters are:
$\begin{split}\epsilon_{\mu}\approx&\sqrt{2\epsilon_{\rm cmb}}\frac{\left(8\sqrt{3}-13\right)\tilde{\mu}^{2}\phi_{0}+\left(4\sqrt{3}-8\right)\tilde{\mu}\phi_{0}^{3}}{2\phi_{0}^{6}}\,,\\ \eta_{\mu}\approx&\frac{(8\sqrt{3}-13)\tilde{\mu}^{2}+(5\sqrt{3}-10)\tilde{\mu}\phi_{0}^{4}}{2\phi_{0}^{6}}\,,\end{split}$ (36)
where we have introduced the rescaled parameter $\tilde{\mu}=\mu/D$ to simplify the expression. The first term in eq.(36) is the cross term between the original and SUSY breaking induced derivatives of the potential in $(V^{\prime})^{2}$. Using $\epsilon_{\rm cmb}=8.88\times 10^{-9}\phi_{0}^{6}$ and requiring $\epsilon_{\mu}<0.05\,\epsilon_{\textrm{cmb}}$, which ensures $\eta_{\mu}\ll\eta_{cmb}$ as well, we get an upper bound on SUSY breaking scale $\mu$:
$\mu<3.4\times 10^{6}\,\left(\frac{\phi_{0}}{M_{\textrm{pl}}}\right)^{6}\,\textrm{GeV}\,.$ (37)
Since we have not found any SUSY particle in collider searches, we conservatively require $\mu>1\textrm{TeV}$. From eq.(37) this implies $\phi_{0}>0.3$, corresponding to $H>10^{8}\ \textrm{GeV}.$
If we increase the SUSY breaking scale while keeping the inflation scale fixed, the Polonyi field will move from the origin to its present minimum at $\sqrt{3}-1$. In Fig. 3 we show the position of the Polonyi field during inflation, by minimizing $V(\phi,z)$ with respect to $z$ for fixed $\phi=\phi_{0}$; here we only consider the real part $z$ of $Z$, fixing the imaginary part to the origin throughout. We see that $z$ tends to stay near the origin when $\tilde{\mu}\ll\phi_{0}^{2}$, and slowly moves to its own vacuum at $\sqrt{3}-1$ as $\tilde{\mu}$ becomes comparable to $\phi_{0}^{2}$.
Figure 3: Position of the Polonyi field $z$ during inflation. Different colors represent different choices of the relative SUSY breaking scale $\tilde{\mu}$. When $\tilde{\mu}\gg\phi_{0}^{2}$, the Polonyi field stays at $\sqrt{3}-1$, whereas for $\tilde{\mu}\ll\phi_{0}^{2}$ the Polonyi field stays close to the origin.
On the other hand, if $W_{\textrm{P}}\gg W_{\textrm{I}}$, the Polonyi field will stay at $Z=\sqrt{3}-1$. In this case, the inflection point conditions $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$ up to $\phi_{0}^{2}$ read:
$\begin{split}V^{\prime}&=0\Rightarrow 8\tilde{B}^{2}+18\sqrt{2}\tilde{B}\tilde{C}\phi_{0}-4\left(\sqrt{3}-2\right)\tilde{B}\tilde{\mu}+18\tilde{C}^{2}\phi_{0}^{2}-3\sqrt{2}\left(\sqrt{3}-3\right)\tilde{C}\tilde{\mu}\phi_{0}+2\tilde{\mu}^{2}=0\,,\\ V^{\prime\prime}&=0\Rightarrow 4\tilde{B}^{2}+18\sqrt{2}\tilde{B}\tilde{C}\phi_{0}-2\left(\sqrt{3}-2\right)\tilde{B}\tilde{\mu}+27\tilde{C}^{2}\phi_{0}^{2}-3\sqrt{2}\left(\sqrt{3}-3\right)\tilde{C}\tilde{\mu}\phi_{0}+\tilde{\mu}^{2}=0\,,\end{split}$ (38)
with the following solutions:
$\tilde{B}\approx 0.4952\tilde{\mu}\,,\quad\tilde{C}\approx-\frac{0.4996\tilde{\mu}}{\phi_{0}}\,.$ (39)
The potential at the inflection point is then:
$V(\phi_{0})\approx 0.1872\,\mu^{2}\phi_{0}^{2}\,.$ (40)
By Taylor expanding around the inflection point and substituting $\tilde{B}\to\tilde{B}(1-\delta B),\phi\to\phi_{0}(1-\delta\phi)$, we find:
$\begin{split}\epsilon&\approx\frac{12.1\left(\delta B+2.44\delta\phi^{2}\right)^{2}}{\phi_{0}^{2}}\,,\\ \eta&\approx\frac{21.6\delta B-24.0\delta\phi}{\phi_{0}^{2}}\,.\end{split}$ (41)
This expansion is similar in structure from the previous cases. Following the same procedure and using $n_{s}=0.9659,\,N_{\textrm{cmb}}=65$, we have:
$\delta B=4.45\times 10^{-6}\phi_{0}^{4}\,,\quad\epsilon_{\textrm{cmb}}=3.92\times 10^{-10}\phi_{0}^{6},\quad\alpha\approx-0.0013\,.$ (42)
We can further deduce the scales of SUSY breaking scale and inflation:
$\mu=\tilde{\mu}D=4.82\times 10^{-8}\phi_{0}^{2}\,,\quad H_{\textrm{inf}}=4.57\times 10^{-9}\phi_{0}^{3}\,,\quad m_{\phi}\approx 2\mu\,.$ (43)
If the Polonyi field already sits in its SUSY breaking minimum during inflation, all relevant energy scales, i.e., Hubble scale $H_{\inf}$, the SUSY breaking scale $\mu$, and inflaton mass $m_{\phi}$, are completely determined by the position of inflection point $\phi_{0}$. The scaling of $H_{\inf}$ and $m_{\phi}$ with $\phi_{0}$ is also as in the non-SUSY version of the model, or as in the SUGRA model without Polonyi sector. The new feature is that $\phi_{0}$, or $H_{\textrm{inf}}$, also determines $\mu$; again demanding $\mu>1$ TeV therefore implies $\phi_{0}>2\times 10^{-4}$ in this set-up. This strong correlation can only be relaxed by lifting the Polonyi field away from the SUSY breaking point $\sqrt{3}-1$.
In Fig. 4, we show how different quantities depend on $\phi_{0}$. For $\phi_{0}\ll 1$ this is described by eqs.(43) and (42), while $\phi_{0}\geq 1$ can again only be treated numerically. As before, we fix the spectral index $n_{s}$ and the number of e-folds $N_{\textrm{cmb}}$, and additionally $\tilde{\mu}$ for better illustration.
The Hubble parameter $H_{\textrm{inf}}$ and the tensor-to-scalar ratio $r$ have the same scaling with $\phi_{0}$ as before. They both increase as $\phi_{0}$ approaches unity, and start to decrease for $\phi_{0}>1$. The running of the spectral index is again almost independent of $\phi_{0}$ and of the order of $10^{-3}$. The scaling of the SUSY breaking scale $\mu$ is rather different. When $\phi_{0}<0.1$, the Polonyi field stays around the SUSY breaking point, and increases with $\phi_{0}$ as eq.(43) suggested. When $\phi_{0}>0.1$, the Polonyi field is shifted away from the SUSY breaking point during inflation. This also leads to a milder increase of $\mu$ along $\phi_{0}$. Once $\phi_{0}$ exceeds $1$, the SUSY breaking scale drops dramatically, which is similar to the behavior of $m_{\phi}$ in the previous case. Requiring $\mu\geq 1$ TeV therefore implies $\phi_{0}\leq 5$ for this value of $\tilde{\mu}$.
(a) Hubble scale $H_{\inf}$ during inflation.
(b) SUSY breaking scale $\mu$.
(c) Tensor to scale ratio $r$.
(d) Running of spectral index $\alpha$.
Figure 4: The dependence of the Hubble parameter during inflation $H_{\textrm{inf}}$ (top left), the SUSY breaking scale $\mu$ (top right), the tensor-to-scalar ratio $r$ (bottom left), and the running of spectral index $\alpha$ (bottom right) on $\phi_{0}$. Different lines represent different choices for the number of e-folds: $N_{\textrm{cmb}}=65$ (blue) and $N_{\textrm{cmb}}=45$ (orange). We fixed $\tilde{\mu}=0.01$, $n_{s}=0.9659$ and $P_{\zeta}=2.1\times 10^{-9}$ in this graph.
So far we fixed $\tilde{\mu}=0.01$. For $\phi_{0}\ll 1$ this choice is in fact irrelevant, since the physical parameters $B,\,\delta B,\,C$ and $\mu$ are all fixed uniquely for given $\phi_{0}$, see eqs.(39), (42) and (43). However, we see in Fig. 5, which shows the relation between $\mu$ and $H_{\inf}$, that this is no longer true for $\phi_{0}\geq 0.1$. Nevertheless, for a given Hubble scale during inflation there will be a maximum SUSY breaking scale it can host, corresponding to the case where SUSY is already broken by the Polonyi field:
$\mu<\frac{4}{\phi_{0}}H_{\inf}\approx 3\times 10^{4}\left(\frac{H_{\inf}}{\textrm{GeV}}\right)^{2/3}\textrm{GeV}\,.$ (44)
Once the Polonyi field moves away from the SUSY breaking point, one will have more freedom to set the SUSY breaking value, depending on the position of the Polonyi field during inflation. If the Polonyi field stays at the origin and only perturbs the potential, the SUSY breaking scale would simply be $\mu=\tilde{\mu}\times D$, where $D$ can be treated as a constant. This explains why the relative ratio of three different cases when they deviate from the straight line in Fig. 5 is almost a constant.
Figure 5: The scale of SUSY breaking $\mu$ vs. the inflationary Hubble scale $H_{\inf}$ on a log-log scale. Different colors represent different choices of relative scale $\tilde{\mu}$. The straight line is the Polonyi field dominated case, where the SUSY breaking scale only depends on the inflection point positions. The right, flat region is an inflation field-dominated region, where the SUSY breaking scale depends linearly on the relative scale $\tilde{\mu}$.
[…]
4 Summary
In this paper we revisited the renormalizable inflection point inflation model in the SUGRA framework. We adopted the minimal assumption that only one canonical field drives inflation. While supersymmetry protects the flatness of the potential from radiative corrections, local SUSY or SUGRA also modifies the potential through non-renormalizable terms. These new terms contribute to slow roll parameters on an equal footing. As in the non-supersymmetric case the shape of the potential is determined by the position $\phi_{0}$ of the inflection point, which is a free parameter of our model. When fixing the well-constrained power spectrum of curvature perturbations and its spectral index, we find $\phi_{0}$ controls the tensor-to-scalar ratio $r$, the running of the spectral index $\alpha$, the Hubble scale during inflation $H_{\textrm{inf}}$, and the physical inflaton mass $m_{\phi}$.
For $\phi_{0}\ll 1$ a perturbative treatment is possible. In this case, $r,\,H_{\textrm{inf}}$ and $m_{\phi}$ are monomial functions of $\phi_{0}$ and reach their maximum around $\phi_{0}\approx 1$. The running of the spectral index $\alpha$ is almost independent of $\phi_{0}$ but depends more strongly on $N_{\textrm{cmb}}$ than in the non-supersymmetric, renormalizable version of the model. The tensor-to-scalar ratio $r$ is always smaller than $10^{-7}$, which is below the sensitivity of any current or planned experiments. The running of the spectral index, which lies in order of $\mathcal{O}(-10^{-3})$, might be probed by the next generation of CMB experiments. The predictions of this SUGRA model are quite similar to those from the renormalizable model. All observables have the same scaling with respect to $\phi_{0}$. Thus, even though the SUGRA potential contains terms up to $\phi^{6}$ while the renormalizable potential only has terms up to $\phi^{4}$, it still provides a relatively reliable estimate of inflationary quantities.
The main difference between the SUGRA case and the renormalizable case appears when $\phi_{0}$ exceeds $1$. In this region the exponential factor $e^{1/2\phi^{2}}$ in the SUGRA case becomes large, which suppresses $r,\,H_{\textrm{inf}}$ and $m_{\phi}$. The energy scales are therefore bounded from above: $H_{\textrm{inf}}<10^{11}\ \textrm{GeV}$ and $m_{\phi}<10^{12}\ \textrm{GeV}$. The renormalizable potential is not able to capture this behavior, which leads to a very different prediction of inflationary observables in this large field scenario.
We further added a SUSY breaking Polonyi sector to the model. If the SUSY breaking scale is much smaller than the Hubble scale, the Polonyi field will stay at the origin and serve as a perturbation to the field. When these two energy scales become comparable, the Polonyi field will move away from the origin and modify the inflation potential. These effects lead to a nontrivial bound between the SUSY breaking scale and the inflation scale. We find that for a TeV scale SUSY breaking, we need the Hubble scale to be larger than $1$ GeV.
It has been pointed out that in the KKLT model, the Hubble scale is always smaller than the gravitino mass $m_{3/2}$ (or SUSY breaking scale $\mu$Kallosh:2004yh . In our model, we find a slightly different conclusion: in some regions, the Hubble scale can be larger than the gravitino mass. In such a scenario supersymmetry will protect the inflaton potential from loop corrections. This should allow larger couplings of the inflaton to Standard Model (super)fields, and thus larger reheat temperature than in the non–supersymmetric version.