### v2

# Inflection Point Inflation in Supergravity

Abstract
In this paper, we study the inflection point inflation generated by a polynomial superpotential and a canonical Kähler potential under the supergravity framework, where only one chiral superfield is needed. We find that the special form of the scalar potential limits the inflationary Hubble parameter to values $\lesssim 10^{10}\,\textrm{GeV}$ and the inflaton mass to $\lesssim 10^{11}\,\textrm{GeV}$. We obtain analytic results for small field cases and present numerical results for large field ones. We find the tensor-to-scalar ratio $r<10^{-8}$ is always suppressed in these models, while the running of spectral index $\alpha\approx\mathcal{O}(-10^{-3})$ may be testable in next-generation CMB experiments. We also discuss the possible effects of a SUSY breaking Polonyi term presented in the superpotential where we find a general upper bound for the SUSY breaking scale for a given value of the Hubble parameter.
1 Introduction
Recent observations of the cosmic microwave background (CMB) by the Planck and BICEP/Keck experiments strongly favor an exponentially expanding period of the early universe Planck:2018jri ; BICEP:2021xfz , which is called inflation Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi . The simplest implementation, slow roll inflation, needs a single scalar “inflaton” field $\phi$ rolling over a flat region in the potential. The flatness of the potential can be parameterized by the following slow roll parameters [^2] We set the reduced Planck Mass $M_{\textrm{pl}}=\sqrt{\frac{1}{8\pi G}}\approx 2.4\times 10^{18}\,\textrm{GeV}=1$ throughout this paper. Lyth:2009imm :
$\epsilon=\frac{1}{2}\left(\frac{V^{\prime}}{V}\right)^{2}\,,\quad\eta=\frac{V^{\prime\prime}}{V}\,,\quad\xi^{2}=\frac{V^{\prime}V^{\prime\prime\prime}}{V^{2}}\,$, (1)
where a prime () denotes the derivative with respect to $\phi$. They are related to the CMB observations Planck:2018jri ; BICEP:2021xfz : [^3] We will comment on the very recent ACT determination ACT:2025fju ; ACT:2025tim of $n_{s}$ in Sec. 3.1, following eq.(31).
$\begin{split}P_{\zeta}&=\frac{V}{24\pi^{2}\epsilon}=(2.1\pm 0.1)\times 10^{-9}\,;\\ n_{s}&=1-6\epsilon+2\eta=0.9659\pm 0.0040\,;\\ r_{0.05}&=16\epsilon<0.036\,;\\ \alpha&=16\epsilon\eta-24\epsilon^{2}-2\xi^{2}=-0.0041\pm 0.0067\,.\end{split}$ (2)
Here $P_{\zeta}$ is the power spectrum of Gaussian curvature perturbations, $n_{s}$ is the spectral index, $r$ is the tensor-to-scalar ratio and $\alpha$ is the running of the spectral index.
Power law potentials $V\propto\phi^{n}$ offer one of the most elegant and simplest realizations of inflation. However, such potentials with $n>1$ have been ruled out by the CMB observations, due to the large tensor-to-scalar ratio predicted by these models. Given this fact, a broad program has been pursued to investigate alternative models for inflation. Among them are the so-called inflection point inflation models, where different $\phi^{n}$ terms collectively provide an inflection point in the potential and accommodate inflation around it. It has been realized using flat directions of the Minimally Supersymmetric Standard Model Allahverdi:2006iq ; Allahverdi:2006cx ; Allahverdi:2006we ; Enqvist:2010vd , in the context of Higgs inflation Okada:2016ssd ; Okada:2017cvy and warped $D$-brane inflation models Baumann:2007np ; Baumann:2007ah . As a generic model, it has been well studied both during and after inflation Hotchkiss:2011am ; Chang:2013cba ; Martin:2013tda ; Martin:2013nzq ; Choi:2016eif ; Dimopoulos:2017xox ; Musoke:2017frr ; Drees:2021wgd ; Drees:2022aea ; Drees:2024hok ; Bernal:2024ykj ; Drees:2025iue .
Supergravity emerges as the low energy limit of superstring theory, which is probably our best bet for a perturbative understanding of quantum gravity. Moreover, supersymmetry protects the (super)potential from loop corrections, which otherwise can easily spoil the required flatness of the inflaton potential. [^4] In large field models, quantum corrections to the Kähler metric can still be dangerous. In this paper we therefore consider inflation models in the supergravity (SUGRA) framework Nanopoulos:1982bv ; Adams:1992bn ; Copeland:1994vg ; Kumekawa:1994gx ; Izawa:1996dv ; Panagiotakopoulos:1997qd ; Linde:1997sj ; Kawasaki:2000yn ; Binetruy:2004hh ; Ellis:2013xoa ; Ellis:2013nxa ; Kallosh:2013yoa ; Carrasco:2015pla . Supergravity models with inflection point inflation have previously been constructed with canonical Kähler potential in Holman:1984yj ; Ross:1995dq ; Mazumdar:2011ih , with logarithmic Kähler potential in Linde:2007jn ; Conlon:2008cj ; Maharana:2015saa ; Gao:2015yha ; Pallis:2023aom , through a stabilizer field in Gao:2016xfv and with extended supergravity in Aldabergenov:2020bpt .
In this paper, we focus on a set–up with a single chiral superfield. A no-go theorem shows that a single chiral superfield with logarithmic Kähler potential can not support inflation Badziak:2008yg ; Ben-Dayan:2008fhy ; Covi:2008cn , unless a modification of Kähler potential is made Ketov:2014qha . Hence, we choose to work with canonical Kähler potential. For simplicity, we consider a polynomial superpotential with three terms. This mirrors non–supersymmetric inflection point inflation with renormalizable potential Drees:2021wgd , and allows us to control the position of the inflection point explicitly. Flexibility and control are among the novel features of our model.
There is a price we have to pay. Like many other inflection point models, parameters in this model are fine-tuned to support inflation and fit the CMB observables. There is no symmetry argument indicating these parameters should take specific values. In this sense, our model belongs to the category where inflation happens accidentally in the parameter space Linde:2007jn . Some progress has been made to reduce the fine-tuning, as in  Enqvist:2010vd ; Hotchkiss:2011am ; Dimopoulos:2017xox . It is beyond the scope of this paper to address this problem.
SUSY must be broken since no superpartner of a Standard Model particle has as yet been detected. Once SUSY breaking terms are included, they may modify the form of the inflaton potential, thereby breaking the slow-roll conditions. These corrections will be under control if the SUSY breaking scale is not too large. We investigate this issue for the simple case where SUSY is broken by the Polonyi model Polonyi:1977pj , which can play an important role in cosmology. In particular, the late time decay of Polonyi particles may adversely impact Big Bang nucleosynthesis (BBN); this is known as the “Polonyi problem” Ellis:1986zt ; Coughlan:1983ci . In F-term inflation models, this leads to a stringent constraint on the Hubble parameter during inflation Ibe:2006am . For D-term inflation, it has been studied in Addazi:2017ulg ; Aldabergenov:2017bjt ; Aldabergenov:2017hvp ; Abe:2018rnu . Here we focus on the impact of the Polonyi superpotential on the inflation potential. It has been shown that no-scale Starobinsky-like inflation is only compatible with a Polonyi sector if the SUSY breaking scale is smaller than $10^{3}\,\textrm{TeV}$ Romao:2017uwa . In this paper, we analyze two limiting cases by comparing the relative strengths of the inflation sector and the SUSY breaking sector. They lead to very different bounds on the SUSY breaking scale, which can be quite large in our model. This is another new result presented in this paper, which defuses the Polonyi problem, since the lifetime of Polonyi particles scales with the inverse third power of their mass; they will thus decay before the onset of BBN if their mass exceeds some tens of TeV.
In this paper we present a systematic study allowing us to scan over the possible parameter space based on the position of the inflection point. We will focus on the similarities and differences between the previously studied non-supersymmetric, renormalizable models and SUGRA models. To be more precise, we consider a SUSY-preserving inflaton field $\phi$, which accommodates a near inflection point in the scalar potential at $\phi=\phi_{0}$. We find that the Hubble parameter and the inflaton mass both increase with $\phi_{0}$ as long as $\phi_{0}<1$. They reach a maximum around $\phi_{0}\sim 1$, then start to decrease because of the exponential factor in the potential. Thus, in such a model, the Hubble parameter can not exceed $\mathcal{O}(10^{10})\ \textrm{GeV}$. The tensor-to-scalar ratio $r$ is much smaller than the current upper bound. The model also predicts a near constant running $\alpha\sim-0.003$, a unique feature that the next generation of observations might be able to test Kohri:2013mxa ; CMB-S4:2016ple ; Munoz:2016owz ; Modak:2021zgb ; Easther:2021rdg ; Bahr-Kalus:2022prj .
The remainder of this paper is organized as follows. In sec. 2 we recap basic results obtained in the non-supersymmetric, renormalizable version of the model and argue why it is interesting to consider this scenario in the SUGRA framework. In sec. 3 we present our model’s analytic and numerical results. We further discuss the SUSY breaking effects in our model. In sec. 4 we summarize the work and draw some conclusions.
2 Non-supersymmetric Inflection Point Model and Beyond
2.1 Potential setup and CMB observables
We start from the most general renormalizable potential of a single real scalar inflaton field $\phi$ as a non-supersymmetric example:
$V(\phi)=b\phi^{2}+c\phi^{3}+d\phi^{4}\,$, (3)
where we have removed the constant and linear terms, so that the minimum of the potential is defined to be at $\phi=0$ with vanishing vacuum energy [^5] A tiny cosmological constant can be added.. Requiring an inflection point, $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$, then leads to:
$b=\frac{9c^{2}}{32d}\,,\quad\phi_{0}=-\frac{3c}{8d}\,.$ (4)
CMB observations indicate the potential is not exactly flat, but rather concave. One way to realize this is to introduce a small deviation from the inflection point conditions in the cubic term. For our purpose it is more convenient to write the coefficients in terms of the inflection point position $\phi_{0}$. The modified potential then reads:
$V(\phi)=d\left(\phi^{4}-\frac{8}{3}\phi_{0}(1-\beta)\phi^{3}+2\phi_{0}^{2}\phi^{2}\right)\,.$ (5)
There are three free parameters in the potential, $d,\phi_{0}$ and $\beta$. In eq.(5), $d$ determines the overall normalization of the potential, which can be matched to the power spectrum of curvature perturbation $P_{\zeta}$ once the other parameters are fixed. The other two parameters govern the shape the potential and hence determine the CMB observables, such as the number of e-folds of inflation after the CMB pivot scale left the horizon, $N_{\textrm{cmb}}$, and the spectral index $n_{s}$. In a “small field” set–up, i.e. for sub-Planckian field values, $\phi_{0}<1$, fixing $N_{\textrm{cmb}}=65$ and eq.(2) require Drees:2021wgd :
$\begin{split}d&=6.61\times 10^{-16}\phi_{0}^{2}\,;\\ \beta&=9.73\times 10^{-7}\phi_{0}^{4}\,.\end{split}$ (6)
This leads to the following predictions:
$\begin{split}b&=1.3\times 10^{-15}\,\phi_{0}^{4}\,;\quad c=1.8\times 10^{-15}\,\phi_{0}^{3}\,;\\ \indent H_{\textrm{inf}}&=8.6\times 10^{-9}\,\phi_{0}^{3}\,;\quad m_{\phi}=5.2\times 10^{-8}\,\phi_{0}^{2}\,;\\ \indent r&=7.1\times 10^{-9}\,\phi_{0}^{6}\,;\quad\alpha=-1.4\times 10^{-3}\,.\end{split}$ (7)
The same ansatz for the potential can also describe “large field” scenarios, where $\phi_{0}>1$. However, in this case no analytical treatment is known. Numerical studies showed that this model can cover the whole allowed parameter space in the $n_{s}-r$ plane and may even lead to double eternal inflation Drees:2022aea .
2.2 Possible Realization in SUGRA and Associated Problem
In SUGRA [^6] For a detailed introduction, please consult Freedman:2012zz . the scalar potential is generated by the real Kähler potential $K$ and the holomorphic Superpotential $W$. Both are supposed to be functions of the complex field $\Phi$, which is the scalar component of a chiral superfield. The Kähler potential also determines the kinetic term of the scalar field (in the convention of ref. Freedman:2012zz ):
$\mathcal{L}_{\textrm{kin}}=-\frac{\partial^{2}K}{\partial\Phi\partial\Phi^{*}}\partial_{\mu}\Phi\partial^{\mu}\Phi^{*}\,.$ (8)
A canonically normalized field thus requires $K=\Phi\Phi^{*}$ and we will use this Kähler potential throughout the paper. We can further define the Kähler covariant derivative as:
$D_{\Phi}W=\frac{\partial W}{\partial\Phi}+\frac{\partial K}{\partial\Phi}W\,=\frac{\partial W}{\partial\Phi}+\Phi^{*}W\,.$ (9)
The $F$-term contribution to the scalar potential then reads:
[…]
Since we assume the inflaton to be a gauge singlet, there is no $D$-term contribution.
For simplicity we choose the superpotential to be a polynomial function of the complex field $\Phi$:
$W(\Phi)=B\Phi^{2}+C\Phi^{3}+D\Phi^{4}\,$, (11)
where we neglect the constant and linear terms as in the renormalizable case. The ansatz (11) ensures that $V(0)=W(0)=0$, i.e. the potential has a supersymmetric stationary point at $\Phi=0$. The three coefficients $B,C,D$ are chosen to be real. The complex $\Phi$ can in general be written as $\Phi=(\phi+i\chi)/\sqrt{2}$, where $\phi$ and $\chi$ are real fields. We want to identify $\phi$ with the inflaton field. By choosing the coefficients in eq.(11) to be real we make sure that the potential $V(\phi,\chi)$ depends only on even powers of $\chi$. Moreover, we make sure that $\partial^{2}V(\phi,\chi)/\partial\chi^{2}>0$ for $\chi=0$ and $\phi\in[0,\phi_{0}]$. We therefore can assume that $\chi=0$ throughout, so that we can ignore this imaginary component when computing the inflationary dynamics. It is not hard to write down the scalar potential $V(\phi)\equiv V(\phi,\chi=0)$ in this setup. We first notice that the two terms in eq.(10) read:
[…]
The combination of them leads to the full potential:
[…]
We first try to match to the non-supersymmetric case by expanding $V$ up to the fourth order:
$V(\phi)\approx 2B^{2}\phi^{2}+3\sqrt{2}BC\phi^{3}+\frac{1}{4}(5B^{2}+9C^{2}+16BD)\phi^{4}\,.$ (14)
Naively these are the leading terms if $\phi\ll 1$. Matching eq.(14) to eq.(3) and using eqs.(6) and (7) implies:
$\begin{split}B&=2.571\times 10^{-8}\phi_{0}^{2}\,,\\ C&=1.615\times 10^{-8}\phi_{0}\times(1-\beta)\,,\\ D&=7.210\times 10^{-10}-8.034\times 10^{-9}\phi_{0}^{2}-5.7\times 10^{-8}(2\beta-\beta^{2})\,.\end{split}$ (15)
Hence for $\phi_{0}\ll 1$ the coefficients of the superpotential would have to scale as $B\propto\phi_{0}^{2},\ C\propto\phi_{0}^{1},\ D\propto\phi_{0}^{0}$. The problem emerges when we substitute these matching conditions back into the full potential eq.(13). Writing the latter as:
$V(\phi)=e^{\frac{1}{2}\phi^{2}}\sum_{n=2}^{10}a_{n}\phi^{n}\,$, (16)
the coefficients $a_{n}$ have the following scaling behavior for $\phi_{0}\ll 1$:
$\begin{split}&a_{2}\propto\phi_{0}^{4};\quad a_{3}\propto\phi_{0}^{3};\quad a_{4}\propto\phi_{0}^{2};\quad a_{5}\propto\phi_{0}^{1};\quad a_{6}\propto\phi_{0}^{0};\\ \indent&a_{7}\propto\phi_{0}^{1};\quad a_{8}\propto\phi_{0}^{0};\quad a_{9}\propto\phi_{0}^{1};\quad a_{10}\propto\phi_{0}^{0}.\end{split}$ (17)
Thus, when we evaluate the value of the potential and its first and second derivatives at $\phi\sim\phi_{0}$, the first five terms (from $\phi^{2}$ to $\phi^{6}$) would contribute with comparable magnitude. Therefore the terms $\propto\phi^{5}$ and $\phi^{6}$ can easily spoil the flatness of the potential, i.e. the expansion only to order $\phi^{4}$ is not self–consistent. For this reason, it is necessary to consider an inflection point model in the full potential eq.(13) (or at least keeping terms up to $\phi^{6}$ in the small field case) rather than trying to directly match the renormalizable, non-supersymmetric potential.
3 Inflection Point Model in SUGRA
3.1 Analytic Analysis of the Model
Requiring that the full potential eq.(13) has an inflection point at $\phi=\phi_{0}$, i.e. $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$, leads to the following solutions [^7] Since we are dealing with high-order polynomial equations there often are several solutions. However, we find that the others usually have $V<0$ at the minimum, leading to a very large and negative cosmological constant.:
$\begin{split}B=&D\frac{\phi_{0}^{2}\left(1152+480\phi_{0}^{2}+72\phi_{0}^{4}+12\phi_{0}^{6}+\phi_{0}^{8}\right)}{2\left(192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}\right)}\,;\\ C=&-D\frac{\sqrt{2}\phi_{0}\left(384+192\phi_{0}^{2}+24\phi_{0}^{4}+8\phi_{0}^{6}+\phi_{0}^{8}\right)}{192+96\phi_{0}^{2}+4\phi_{0}^{6}+\phi_{0}^{8}}\,.\end{split}$ (18)
The inflection point condition allows us to bypass the $\eta$ problem for the canonical Kähler potential Copeland:1994vg ; Stewart:1994ts , which assumes that a single term in the second derivatives $V^{\prime\prime}$, arising from the exponential factor, dominates over the other possible ones. For a scalar potential accommodating an inflection point, this assumption doesn’t hold, and a flat potential appears only when we go beyond it. As already noted in the Introduction, this requires tuning of a priori unrelated parameters, in particular of the ratio $B/C$, as described by eqs.(18).
For $\phi_{0}\ll 1$ we can approximate the above solutions by their leading order results:
$B\approx D\times 3\phi_{0}^{2}\,,\quad C\approx D\times(-2\sqrt{2}\phi_{0})\,.$
The full potential at $\phi\leq\phi_{0}$ can be further simplified if we only include terms up to $\phi_{0}^{6}$:
$V(\phi)\approx 2D^{2}\phi^{2}(\phi^{2}-3\phi\phi_{0}+3\phi_{0}^{2})^{2}\,.$ (19)
This expansion now is self-consistent, i.e. the higher order terms, starting at ${\cal O}(\phi^{7})$, are indeed suppressed. At the inflection point $\phi_{0}$, the potential reads:
$V(\phi_{0})\approx 2D^{2}\phi_{0}^{6}\,.$ (20)
The potential (19) has a minimum at $\phi=0$, with $V(0)=0$. The algebraic equation $D_{\phi}W\pm\sqrt{3}W=0$ has no real root besides $\phi=0$ when using eq.(18), which can be verified by checking the discriminant of the equation with additional combination of coefficients Rees:1922 . Hence the full potential of eq.(13) is positive semi-definite, i.e. $V(\phi)\geq 0\ \forall\phi$. [This is immediately obvious for the simplified potential of eq.(19).] As in the non-supersymmetric case, in the small field scenario $\phi_{0}\ll 1$ we can get semi-analytic results for inflationary observables by expanding the slow-roll parameter around the inflection point via the ansatz $\phi=\phi_{0}(1-\delta\phi)$, and adding a deviation from the strict inflection point condition via $B\to B+D\times\delta B$:
$\begin{split}\epsilon&=\frac{1}{2}\left(\frac{V^{\prime}}{V}\right)^{2}=\frac{1}{2}\left(\frac{2\delta B+6\phi_{0}^{2}\delta\phi^{2}}{\phi_{0}^{3}}\right)^{2}\,,\\ \eta&=\frac{V^{\prime\prime}}{V}=\frac{10.5\,\delta B-24\,\delta\phi}{2\phi_{0}^{2}}\,,\\ \xi^{2}&=\frac{V^{\prime}V^{\prime\prime\prime}}{V^{2}}=\frac{24\delta B+72\phi_{0}^{2}\delta\phi^{2}}{\phi_{0}^{6}}\,.\end{split}$ (21)
We only keep terms up to linear order in $\delta B$ and quadratic in $\delta\phi$; this will turn out to be sufficient.
We are now ready to discuss which $\delta\phi$ and $\delta B$ reproduce the CMB observations. In our model the duration of inflation is controlled by $\eta$ since $\epsilon\ll|\eta|$. The beginning (when the pivot scale $k_{*}=0.05\,\textrm{Mpc}^{-1}$ crossed out of the horizon) and end of observable inflation are given by:
$\eta_{\textrm{cmb}}=\frac{n_{s}-1}{2}\,,\quad\eta_{\textrm{end}}=-1\,$, (22)
Solving the second eq.(21) for $\delta\phi$, we get:
$\delta\phi=-\frac{2\phi_{0}^{2}\eta-10.5\delta B}{24}\,.$ (23)
The number of e-folds $N_{\textrm{cmb}}$ is given by:
$\begin{split}N_{\textrm{cmb}}&=\int_{\phi_{\textrm{cmb}}}^{\phi_{\textrm{end}}}\frac{1}{\sqrt{2\epsilon}}d\phi\\ \indent&=\int_{\phi_{\textrm{cmb}}}^{\phi_{\textrm{end}}}\frac{\phi_{0}^{3}}{2\delta B+6\phi_{0}^{2}\delta\phi^{2}}d\phi\\ \indent&=-\int_{\delta\phi_{\textrm{cmb}}}^{\delta\phi_{\textrm{end}}}\frac{\phi_{0}^{4}}{2\delta B+6\phi_{0}^{2}\delta\phi^{2}}d\delta\phi\\ \indent&=\frac{\phi_{0}^{2}}{6\sqrt{\beta}}\left[\arctan\left({\frac{\delta\phi_{\textrm{end}}}{\sqrt{\beta}}}\right)-\arctan\left({\frac{\delta\phi_{\textrm{cmb}}}{\sqrt{\beta}}}\right)\right]\,.\end{split}$ (24)
In the last step we have switched from the absolute deviation $\delta B$ to the relative one $\beta=D\times\delta B/B=\delta B/(3\phi_{0}^{2})$ in order to factor out $\phi_{0}^{2}$ in the denominator; except for the first line, eq.(24) also only holds for $\phi_{0}\ll 1$. Recall that the arc-tangent functions can be at most $\pi/2$. Numerically, requiring $N_{\textrm{cmb}}=50$ yields $\beta=2.7\times 10^{-5}\phi_{0}^{4}$ and $\delta B=8.2\times 10^{-5}\phi_{0}^{6}$. From the second eq.(2) and remembering $|\epsilon|\ll|\eta|$ we see that $|\eta|\geq 0.015$ when and after CMB scales crossed out of the horizon. Hence we can neglect the second term in eq.(23) for $\phi_{0}\ll 1$, and determine the inflation period by:
$\delta\phi=-\frac{\phi_{0}^{2}}{12}\eta\,.$ (25)
Inserting this into the last line of eq.(24) yields
$N_{\textrm{cmb}}=\frac{\phi_{0}^{2}}{6\sqrt{\beta}}\left[\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{end}}}{12\sqrt{\beta}}}\right)-\arctan\left({\frac{-\phi_{0}^{2}\eta_{\textrm{cmb}}}{12\sqrt{\beta}}}\right)\right]\,$, (26)
which can be solved numerically. For example, setting $n_{s}=0.9659\,,N_{\textrm{cmb}}=45$ would result in $\delta B=6.1\times 10^{-5}\phi_{0}^{6}$.
Having fixed $\delta B$ and the initial $\delta\phi$, the overall scale of inflation, and hence $D$, is determined by the power of Gaussian curvature perturbations. To this end we first calculate $\epsilon$ at $\phi_{\textrm{cmb}}$:
$\epsilon_{\textrm{cmb}}=8.88\,\times 10^{-9}\phi_{0}^{6}\,$, (27)
Using eqs.(2) and (20) we find that the normalization factor is independent of $\phi_{0}$:
$D^{-2}=\frac{2\phi_{0}^{6}}{P_{\zeta}24\pi^{2}\epsilon_{\textrm{cmb}}}=4.52\times 10^{14}\,.$ (28)
It is then straightforward to evaluate the value of the Hubble parameter during inflation and the physical mass of the inflaton after inflation:
$H_{\textrm{inf}}=\sqrt{\frac{V(\phi_{0})}{3}}=3.84\times 10^{-8}\phi_{0}^{3}\,,\quad m_{\phi}=\sqrt{4B^{2}}=2.81\times 10^{-7}\phi_{0}^{2}\,.$ (29)
The running of the spectral index $\alpha$ can also be determined:
$\alpha=16\epsilon\eta-24\epsilon^{2}-2\xi^{2}\approx-2\xi^{2}=-2\frac{24\delta B+72\phi_{0}^{2}\delta\phi^{2}}{\phi_{0}^{6}}\approx-0.0030\,.$ (30)
This running of the spectral index is a feature of our model, since it does not depend on $\phi_{0}$.
However, our numerical results do depend on $N_{\textrm{cmb}}$. Taking as second example the rather large value $N_{\textrm{cmb}}=65$ while keeping $n_{s}=0.9659$, we get:
$\begin{split}\delta B=2.60\times 10^{-5}\phi_{0}^{6}\,&,\quad\epsilon_{\textrm{cmb}}=1.59\times 10^{-9}\phi_{0}^{6}\,,\\ D^{-2}=2.53\times 10^{15}\,&,\quad H_{\textrm{inf}}=1.62\times 10^{-8}\phi_{0}^{3}\,,\\ m_{\phi}=1.19\times 10^{-7}\phi_{0}^{2}\,&,\quad\alpha=-0.0015\,.\end{split}$ (31)
The tensor-to-scalar ratio $r=16\epsilon$ is always too small to reach the sensitivity of any currently conceivable observation. However, the S4CMB experiment, together with small-scale structure information (e.g. on the Lyman-$\alpha$ forest) could achieve $10^{-3}$ sensitivity for $\alpha$, which would test our model Kohri:2013mxa ; CMB-S4:2016ple ; Munoz:2016owz ; Modak:2021zgb ; Easther:2021rdg ; Bahr-Kalus:2022prj .
Recently, the ACT experiment measured a slightly larger spectral index when combining their results with DESI and Planck data, $n_{s}=0.9743\pm 0.0034$ DESI:2024mwx ; ACT:2025fju ; ACT:2025tim . Using this central value gives us slightly different predictions. Taking $N_{\textrm{cmb}}=45$ as an example, we have:
$\begin{split}\delta B=6.88\times 10^{-5}\phi_{0}^{6}\,&,\quad\epsilon_{\textrm{cmb}}=1.05\times 10^{-8}\phi_{0}^{6}\,,\\ D^{-2}=3.85\times 10^{14}\,&,\quad H_{\textrm{inf}}=4.16\times 10^{-8}\phi_{0}^{3}\,,\\ m_{\phi}=3.06\times 10^{-7}\phi_{0}^{2}\,&,\quad\alpha=-0.0035\,.\end{split}$ (32)
while taking $N_{\textrm{cmb}}=65$, we have:
$\begin{split}\delta B=2.83\times 10^{-5}\phi_{0}^{6}\,&,\quad\epsilon_{\textrm{cmb}}=2.02\times 10^{-9}\phi_{0}^{6}\,,\\ D^{-2}=1.99\times 10^{15}\,&,\quad H_{\textrm{inf}}=1.83\times 10^{-8}\phi_{0}^{3}\,,\\ m_{\phi}=1.34\times 10^{-7}\phi_{0}^{2}\,&,\quad\alpha=-0.0015\,.\end{split}$ (33)
One can see a shift in the spectral index leaves a subtle change on inflationary observables. The larger value of the spectral index from ACT would prolong the duration of inflation in the field space. Hence for a fixed number of e-folds $N_{\textrm{cmb}}$, ACT results will predict larger $\epsilon_{\textrm{cmb}}$, larger Hubble scale value $H_{\textrm{inf}}$ and inflaton mass $m_{\phi}$. This effect can be clearly seen by comparing the results in eq.(31) and (33). However, the differences between the model parameters derived from PLANCK and ACT observations with fixed $N_{\rm CMB}$ are much smaller than the variations of these parameters when $N_{\rm CMB}$ is varied between $45$ and $65$. We will therefore continue to use the PLANCK derived parameter values and estimate their uncertainty by considering these two values of $N_{\rm CMB}$.
Before moving to the next part, we would like to briefly mention the possibility of Ultra-slow-roll (USR) in our model, as inflection point inflation is a well-known example to study the USR process. To have a USR period, we need the kinetic energy of the inflation field to be much larger than its slow-roll kinetic energy Dimopoulos:2017ged . In order to have a stable USR, the second order derivative of the potential has to be positive: $V^{\prime\prime}>0$ Pattison:2018bct . The first condition is up to the initial conditions of the inflaton field dynamics. In our calculation, we have assumed the evolution starts from the slow-roll phase. This assumption can be relaxed to incorporate USR. The second condition is stricter. Throughout the observable inflation, our potential features a negative slow roll parameter $\eta$, i.e., $V^{\prime\prime}<0$. This excludes a stable USR phase.
By comparing with the predictions of the renormalizable model given in eqs.(7), we see that $H_{\textrm{inf}}$ and $m_{\phi}$ scale with $\phi_{0}$ in the same way in both cases, with roughly a factor $2$ difference in coefficients. Even though the SUGRA potential is more complicated and scales as $\phi^{6}$ in the simplest limit, they thus make very similar predictions for the inflationary observables. In particular, the overall coefficient $D$ of the sixth order SUGRA potential is independent of $\phi_{0}$, while in the non-SUSY case the coefficient of the quartic potential $d\propto\phi_{0}^{2}$; hence in both cases $V(\phi\simeq\phi_{0})\propto\phi_{0}^{6}$.
We also note that the curvature of the potential is negative for an extended range of $\phi$ below $\phi_{0}$. In the non-SUSY case, this holds for $\phi/\phi_{0}\in[1/3,1]$, independently of the value of $\phi_{0}$; in the SUGRA case with $\phi_{0}\ll 1$, this region extends to $\phi/\phi_{0}\in[1/4,1]$. The minimum of the curvature occurs at $\phi\simeq 0.54\phi_{0}$, closer to the origin than in the non-SUSY case, with a value just below $0.23~m_{\phi}^{2}$, smaller in magnitude than in the non-SUSY case. The latter reduces the tachyonic instability while the former increases it. Hence, we expect the non-perturbative effects after inflation studied in Drees:2025iue for the non-SUSY case would be similar in the SUGRA version of the model.
When $\phi_{0}$ is larger than unity, it is hard to make a comprehensive analytic analysis, but we can still understand the model qualitatively. To this end we first formally rewrite the potential as:
$V(\phi)=e^{\phi^{2}/2}P(\phi)\,.$ (34)
The slow roll parameter then scales like $\eta\propto\phi^{2}\times f(\delta\phi)$ with $f(0)\approx 0$. The duration of inflation is still controlled by $\delta\eta=\eta_{\textrm{cmb}}-\eta_{\textrm{end}}\approx 1$, the larger $\phi_{0}$, the smaller $\delta\phi$ should be. The resulting decrease in the integration range in eq.(24) has to be compensated by increasing the integrand, by a reduction of $\epsilon$. Since the ratio between the potential $V(\phi_{0})$ and $\epsilon$ is fixed by the power of curvature perturbations, an increase of $\phi_{0}$ would eventually lead to a decrease of the potential and hence of the Hubble parameter. Moreover, the potential near $\phi_{0}$ is increased by the exponential factor, which becomes unity at the origin. Hence, the mass of inflaton should be suppressed by $e^{-\frac{1}{2}\phi_{0}^{2}}$.
As a brief summary, by fixing $n_{s}$ and $N_{\textrm{cmb}}$, we find the inflection point model would always give a tiny tensor-to-scalar ratio $r$ and a constant running of spectral index $\alpha$. The Hubble scale of inflation first increases with increasing inflection point position $\phi_{0}$, and then decreases once $\phi_{0}$ exceeds $1$. The inflaton mass follows the same pattern but drops much faster in the second phase. In the next section, we confirm these expectations by showing some numerical results.
3.2 Numerical Results of the Model
In this section, we present our numerical results. We introduce three steps to scan the allowed parameter space of our model, following the same spirit as our analytic treatment:
  • We choose $\phi_{0}$ as a free parameter and solve the inflection point equations $V^{\prime}=V^{\prime\prime}=0$ to find corresponding values of $B/D$ and $C/D$. We pick the solution that will generate a positive semi-definite potential, see eqs.(18).
  • We slightly deform the potential by $B\to B+\delta B$. CMB scales start to leave the horizon at $\phi_{\textrm{cmb}}$, which is determined by $1+2\eta=n_{s}$ since still $\epsilon\ll|\eta|$ in all cases. Inflation ends at $\phi_{\textrm{end}}$, which is determined by $\eta=-1$. We find that both the start point and the end point mildly depend on $\delta B$. The correct $\delta B$ is given by fixing $N_{\textrm{cmb}}$, for which we consider the range from $45$ to $65$.
  • Having fixed $\delta B$, we can recalculate the slow roll parameters at the pivot scale $\eta_{\rm cmb}$ and $\epsilon_{\rm cmb}$. We then determine the correct normalization $D$ of the potential by requiring $P_{\zeta}=2.1\times 10^{-9}$. This allows us to compute the Hubble value and the inflaton mass.
Figure 1: Rescaled inflation potential for different choices of the location of the inflection point $\phi_{0}$. Here $V_{0}=V(\phi_{0})$ is the value of the potential at the inflection point. Blue, orange, green, and red curves corresponding to $\phi_{0}=0.1,\,1,\,3$ and $5$, respectively.
We begin by showing four inflaton potentials with different choices of $\phi_{0}$ in Fig. 1. For comparison, these potentials are rescaled by their values at the inflection point. When $\phi_{0}<1$, the shape of the potential becomes independent of $\phi_{0}$ for $\phi\leq\phi_{0}$. Increasing $\phi_{0}$ beyond $1$ shortens the flat plateau; it also makes it even flatter, which is difficult to see in this figure.
(a) Hubble scale $H_{\inf}$ during inflation.
(b) Inflaton mass $m_{\phi}$.
(c) Tensor to scale ratio $r$.
(d) Running of spectral index $\alpha$.
Figure 2: The dependence of the Hubble parameter $H_{\inf}$ during inflation, the inflaton mass $m_{\phi}$, tensor-to-scalar ratio $r$, and the running of spectral index $\alpha$ on the position $\phi_{0}$ of the inflection point. Different lines represent different choices of the number of e-folds: $N_{\textrm{cmb}}=65$ (blue) and $N_{\textrm{cmb}}=45$ (orange). We fixed $n_{s}=0.9659$ and $P_{\zeta}=2.1\times 10^{-9}$.
The corresponding values of the Hubble parameter during inflation can be found in the top left frame of Fig. 2, where we use blue and orange lines to represent different values of $N_{\textrm{cmb}}$. As expected the Hubble scale first increases with increasing $\phi_{0}$, then drops once $\phi_{0}>1$. There is no lower bound on $H_{\rm inf}$ from the pure model perspective. However, the maximum value is determined by the special shape of the potential and can never exceed $10^{11}\ \textrm{GeV}$. It obeys a power law when $\phi_{0}$ is small, which agrees with our analytic estimation.
The relation between inflaton mass $m_{\phi}$ and inflection point $\phi_{0}$, shown in the top right frame, follows a similar pattern as the Hubble scale when $\phi_{0}$ is small, albeit with different power. However, the inflaton mass drops dramatically for $\phi_{0}>1$, due to the exponential suppression discussed at the end of the previous section. When $\phi_{0}\approx 10$, the inflaton mass could be as low as $1\ \textrm{GeV}$. This tiny value differs from the Hubble scale by more than nine orders of magnitude. Thus our model offers a way to separate the Hubble and inflaton mass scales.
The running of the spectral index $\alpha$ is shown in the bottom right frame. It remains independent of $\phi_{0}$ even for $\phi_{0}\geq 1$. In contrast to the non-SUSY version of the model, $\alpha$ strongly depends on $N_{\textrm{cmb}}$. When $N_{\textrm{cmb}}=45$, $\alpha\approx-0.0032$. Increasing $N_{\textrm{cmb}}$ reduces the absolute value of $\alpha$, reaching $\alpha\approx-0.0013$ for $N_{\textrm{cmb}}=65$.
Finally, the bottom left frame of Fig. 2 shows the relationship between the tensor-to-scalar ratio $r$ and $\phi_{0}$. It follows a similar pattern as the Hubble scale. However, since we find $r<10^{-7}$, a positive detection of tensor modes by current or near-future experiments would exclude our model.
In principle one can thus distinguish the small and large field scenarios through the tensor-to-scalar ratio. If we restrict the inflection point position so that it should not be much larger than the Planck scale, $\phi_{0}<10$, then a small tensor to scalar ratio $r<10^{-10}$ would favor the small field scenario. If the tensor-to-scalar ratio is between $10^{-8}$ and $10^{-10}$, we need to check further the inflation mass. It cannot be measured directly; however, generally speaking, a large reheating temperature requires a large inflation mass, which can only be realized in the small field version of our model. The reheating dynamics can in principle be probed via the observation of primordial high-frequency gravitational waves Bernal:2023wus ; Xu:2024fjl ; Jiang:2024akb ; Xu:2025wjq . Unfortunately none of these observations are feasible with near–future technology.
3.3 SUSY Breaking by a Polonyi Field
Clearly SUSY must be broken in any realistic model. In this subsection we investigate if the existence of a SUSY breaking sector would change the inflation potential significantly. For simplicity we consider the classical Polonyi ansatz Polonyi:1977pj , where a single chiral superfield $Z$ with a linear superpotential is introduced to break SUSY:
[…]
where $\mu$ essentially sets the SUSY breaking scale; we explicitly include a factor $M_{\textrm{pl}}$ here to ensure the correct dimension of $\mu$. Both $\Phi$ and $Z$ are complex fields. As before, we want the real part of $\Phi$ to be the inflaton field $\phi$. $Z$ is the Polonyi field whose vacuum expectation value $\langle Z\rangle$ is the only source of SUSY breaking after inflation, when $\langle\phi\rangle=0$. SUSY breaking with vanishing vacuum energy requires $\beta_{\rm P}=2-\sqrt{3}$, which gives the gravitino mass $m_{3/2}=\mu e^{2-\sqrt{3}}$ when $Z$ stays at the SUSY breaking minimum at $Z=\langle Z\rangle=(\sqrt{3}-1)M_{\textrm{pl}}$.
Let’s first consider the case where the Polonyi sector gives a small perturbation to the inflation potential. For $\phi_{0}<1$ our previous results suggest:
$\begin{split}\tilde{B}&=\frac{B}{D}\approx 3\phi_{0}^{2}\,,\\ \tilde{C}&=\frac{C}{D}\approx-2\sqrt{2}\phi_{0}\,,\\ D&\approx 4.7\times 10^{-8}\,.\end{split}$ (36)
We require that the Polonyi field does not change the inflation potential significantly, which means $|W_{\textrm{I}}|\gg|W_{\textrm{P}}|$ during inflation. The existence of the SUSY breaking term will not alter the slow roll parameters significantly if
$\epsilon_{\mu}\ll\epsilon_{\rm cmb}\quad{\rm and}\quad\eta_{\mu}\ll\eta_{\rm cmb}\,$, (37)
where the subscript $\mu$ means the additional contribution to the slow roll parameter due to the SUSY breaking term.
We assume that during inflation the Polonyi field stays at the origin, $Z=0$, which will be verified later. Under this assumption, the additional contribution to the inflaton potential reads:
$V_{\mu}(\phi)=\frac{{\rm e}^{\frac{1}{2}\phi^{2}}}{4}\mu\left[(-4\beta_{\rm P}B\phi^{2}+2\beta_{\rm P}(B+D)\phi^{4}+\sqrt{2}\beta_{\rm P}C\phi^{5}+\beta_{\rm P}D\phi^{6})+\mu(4-12\beta_{\rm P}^{2}+2\beta_{\rm P}^{2}\phi^{2})\right]\,.$ (38)
After substituting $\beta_{\rm P}=2-\sqrt{3}$, the additional contributions to slow roll parameters are:
[…]
where we have introduced the rescaled parameter $\tilde{\mu}=\mu/D$ to simplify the expression. The first term in eq.(39) is the cross term between the original and SUSY breaking induced derivatives of the potential in $(V^{\prime})^{2}$. Using $\epsilon_{\rm cmb}=8.88\times 10^{-9}\phi_{0}^{6}$ and requiring $\epsilon_{\mu}<0.05\,\epsilon_{\textrm{cmb}}$, which ensures $\eta_{\mu}\ll\eta_{cmb}$ as well, we get an upper bound on SUSY breaking scale $\mu$:
$\mu<3.4\times 10^{6}\,\left(\frac{\phi_{0}}{M_{\textrm{pl}}}\right)^{6}\,\textrm{GeV}\,.$ (40)
Since we have not found any SUSY particle in collider searches, we conservatively require $\mu>1\ \textrm{TeV}$. From eq.(40) this implies $\phi_{0}>0.3$, corresponding to $H>10^{8}\ \textrm{GeV}.$
If we increase the SUSY breaking scale while keeping the inflation scale fixed, the Polonyi field will move from the origin to its present minimum at $\sqrt{3}-1$. In Fig. 3 we show the position of the Polonyi field during inflation, by minimizing $V(\phi,z)$ with respect to $z$ for fixed $\phi=\phi_{0}$; here we only consider the real part $z$ of $Z$, fixing the imaginary part to the origin throughout. We see that $z$ tends to stay near the origin when $\tilde{\mu}\ll\phi_{0}^{2}$, and slowly moves to its own vacuum at $\sqrt{3}-1$ as $\tilde{\mu}$ becomes comparable to $\phi_{0}^{2}$.
Figure 3: Position of the Polonyi field $Z$ during inflation. Different colors represent different choices of the relative SUSY breaking scale $\tilde{\mu}$. When $\tilde{\mu}\gg\phi_{0}^{2}$, the Polonyi field stays at $\sqrt{3}-1$, whereas for $\tilde{\mu}\ll\phi_{0}^{2}$ the Polonyi field stays close to the origin.
On the other hand, if $|W_{\textrm{P}}|\gg|W_{\textrm{I}}|$, the Polonyi field will stay at $Z=\sqrt{3}-1$; this requires $|\mu|\gg|D|\phi_{0}^{4}$. Focusing on small field solutions, and imposing the slightly stronger condition $|\mu|\gg|D|\phi_{0}^{2}$, the inflection point conditions $V^{\prime}(\phi_{0})=V^{\prime\prime}(\phi_{0})=0$ up to terms ${\cal O}(\phi_{0}^{2})$ read:
[…]
with the following solutions: [^8] This solution implies $B^{2}\ll C^{2}$ as long as $\phi_{0}\ll 1$, which a posteriori justifies dropping the quartic term $\propto B^{2}$ in eq.(13). Moreover, eqs.(42) and our assumption $|\tilde{\mu}|\gg\phi_{0}^{2}$ imply that the third quartic term in eq.(13), $\propto BD$, is also negligible. We note that without this additional assumption, higher than quartic powers of $\phi$ need to be considered, in which case no simple analytical treatment is possible.
$\tilde{B}\approx 0.4952\tilde{\mu}\,,\quad\tilde{C}\approx-\frac{0.4996\tilde{\mu}}{\phi_{0}}\,$, (42)
i.e. the ratio $B/C$ again needs to be tuned for inflection point inflation to occur. The potential at the inflection point is then:
$V(\phi_{0})\approx 0.1872\,\mu^{2}\phi_{0}^{2}\,.$ (43)
By Taylor expanding around the inflection point and substituting $\tilde{B}\to\tilde{B}(1-\delta B),\phi\to\phi_{0}(1-\delta\phi)$, we find:
$\begin{split}\epsilon&\approx\frac{12.1\left(\delta B+2.44\delta\phi^{2}\right)^{2}}{\phi_{0}^{2}}\,,\\ \eta&\approx\frac{21.6\delta B-24.0\delta\phi}{\phi_{0}^{2}}\,.\end{split}$ (44)
This expansion is similar in structure to the previous cases. Following the same procedure and using $n_{s}=0.9659,\,N_{\textrm{cmb}}=65$, we have:
$\delta B=4.45\times 10^{-6}\phi_{0}^{4}\,,\quad\epsilon_{\textrm{cmb}}=3.92\times 10^{-10}\phi_{0}^{6},\quad\alpha\approx-0.0013\,.$ (45)
We can further deduce the scales of SUSY breaking scale and inflation:
$\mu=\tilde{\mu}D=4.82\times 10^{-8}\phi_{0}^{2}\,,\quad H_{\textrm{inf}}=4.57\times 10^{-9}\phi_{0}^{3}\,,\quad m_{\phi}\approx 2\mu\,.$ (46)
If the Polonyi field already sits in its SUSY breaking minimum during inflation, all relevant energy scales, i.e., Hubble scale $H_{\inf}$, the SUSY breaking scale $\mu$, and inflaton mass $m_{\phi}$, are completely determined by the position of inflection point $\phi_{0}$. The scaling of $H_{\inf}$ and $m_{\phi}$ with $\phi_{0}$ is again as in the non-SUSY version of the model, or as in the SUGRA model without Polonyi sector. The new feature is that $\phi_{0}$, or $H_{\textrm{inf}}$, also determines $\mu$; again demanding $\mu>1$ TeV therefore implies $\phi_{0}>2\times 10^{-4}$ in this set-up. This strong correlation can only be relaxed by lifting the Polonyi field away from the SUSY breaking point $\sqrt{3}-1$.
(a) Hubble scale $H_{\inf}$ during inflation.
(b) SUSY breaking scale $\mu$.
(c) Tensor to scale ratio $r$.
(d) Running of spectral index $\alpha$.
Figure 4: The dependence of the Hubble parameter during inflation $H_{\textrm{inf}}$ (top left), the SUSY breaking scale $\mu$ (top right), the tensor-to-scalar ratio $r$ (bottom left), and the running of spectral index $\alpha$ (bottom right) on $\phi_{0}$. Different lines represent different choices for the number of e-folds: $N_{\textrm{cmb}}=65$ (blue) and $N_{\textrm{cmb}}=45$ (orange). We fixed $\tilde{\mu}=0.01$, $n_{s}=0.9659$ and $P_{\zeta}=2.1\times 10^{-9}$ in this graph.
In Fig. 4, we show how different quantities depend on $\phi_{0}$. For $\phi_{0}\ll 1$ and $\phi_{0}^{2}\ll\tilde{\mu}$ this is described by eqs.(46) and (45). For our choice $\tilde{\mu}=0.01$ the latter condition is stronger; indeed, we observe some deviations from eqs.(46) once $\phi_{0}\gtrsim 0.1$. As before, we fix the spectral index $n_{s}$ and the number of e-folds $N_{\textrm{cmb}}$.
The Hubble parameter $H_{\textrm{inf}}$ and the tensor-to-scalar ratio $r$ have the same scaling with $\phi_{0}$ as before. They both increase as $\phi_{0}$ approaches unity, and start to decrease for $\phi_{0}>1$. The running of the spectral index is again almost independent of $\phi_{0}$ and of the order of $10^{-3}$. The scaling of the SUSY breaking scale $\mu$ is rather different. When $\phi_{0}<0.1$, the Polonyi field stays around the SUSY breaking point, and increases with $\phi_{0}$ as eq.(46) suggested. When $\phi_{0}>0.1$, the Polonyi field is shifted away from the SUSY breaking point during inflation. This also leads to a milder increase of $\mu$ along $\phi_{0}$. Once $\phi_{0}$ exceeds $1$, the SUSY breaking scale drops dramatically, which is similar to the behavior of $m_{\phi}$ in the previous case. Requiring $\mu\geq 1$ TeV therefore implies $\phi_{0}\leq 5$ for this value of $\tilde{\mu}$.
Figure 5: The scale of SUSY breaking $\mu$ vs. the inflationary Hubble scale $H_{\inf}$ on a log-log scale. Different colors represent different choices of relative scale $\tilde{\mu}$. The straight line is the Polonyi field dominated case, where the SUSY breaking scale only depends on the inflection point positions. The right, flat region is an inflaton field dominated region, where the SUSY breaking scale depends linearly on the relative scale $\tilde{\mu}$.
So far we fixed $\tilde{\mu}=0.01$. For $\phi_{0}^{2}\ll\tilde{\mu}$ this choice is in fact irrelevant, since the physical parameters $B,\,\delta B,\,C$ and $\mu$ are all fixed uniquely for given $\phi_{0}$, see eqs.(42), (45) and (46). However, we see in Fig. 5, which shows the relation between $\mu$ and $H_{\inf}$, that this is no longer true for $\phi_{0}\geq\sqrt{\tilde{\mu}}$. Nevertheless, this figure shows that for a given Hubble scale during inflation there will be a maximum SUSY breaking scale it can host, given by the analytical solution of eq.(46):
$\mu<\frac{10.5}{\phi_{0}}H_{\inf}\approx 2.3\times 10^{4}\left(\frac{H_{\inf}}{\textrm{GeV}}\right)^{2/3}\textrm{GeV}\,.$ (47)
Recall that this corresponds to a scenario where the Polonyi field remains at its current value throughout inflation. If it moves away from this value during inflation, one will have more freedom to set the value of $\mu$, depending on the position of the Polonyi field during inflation. In the opposite extreme, where the Polonyi field stays at the origin during inflation and only perturbs the inflaton potential, the SUSY breaking scale would simply be $\mu=\tilde{\mu}\times D$, where $D$ can be treated as a constant. This explains why the relative ratio of three different cases when they deviate from the straight line in Fig. 5 is almost a constant.
We conclude that only the light cyan region below the topmost line in Fig. 5 is accessible in our model. Different choices of $\tilde{\mu}$ will leave the straight line at different Hubble scales, and can thus populate this region, always keeping in mind that $\mu>1$ TeV is needed for phenomenological reasons. The resulting lower bound on the Hubble scale during inflation is around $1~\textrm{GeV}$. This bound is much smaller than the naive estimate of $10^{8}$ GeV we derived below eq.(40) from the requirement that the Polonyi sector can be treated as a small perturbation.
4 Summary
In this paper we revisited the inflection point inflation model in the SUGRA framework. We adopted the minimal assumption that only one canonical field drives inflation. While supersymmetry protects the flatness of the potential from radiative corrections, local SUSY or SUGRA also modifies the potential through non-renormalizable terms. These new terms contribute to slow roll parameters on an equal footing. As in the non-supersymmetric case the shape of the potential is determined by the position $\phi_{0}$ of the inflection point, which is a free parameter of our model. When fixing the well-constrained power spectrum of curvature perturbations and its spectral index, we find $\phi_{0}$ controls the tensor-to-scalar ratio $r$, the Hubble scale during inflation $H_{\textrm{inf}}$, and the physical inflaton mass $m_{\phi}$.
For $\phi_{0}\ll 1$ a perturbative treatment is possible. In this case, $r,\,H_{\textrm{inf}}$ and $m_{\phi}$ are monomial functions of $\phi_{0}$ and reach their maximum around $\phi_{0}\approx 1$. The running of the spectral index $\alpha$ is almost independent of $\phi_{0}$ but depends more strongly on $N_{\textrm{cmb}}$ than in the non-supersymmetric version of the model. The tensor-to-scalar ratio $r$ is always smaller than $10^{-7}$, which is below the sensitivity of any current or planned experiments. The running of the spectral index, which is $\mathcal{O}(-10^{-3})$, might be probed by the next generation of CMB experiments. The predictions of this SUGRA model are quite similar to those from the non-supersymmetric model. All observables have the same scaling with respect to $\phi_{0}$. Thus, even though the SUGRA potential contains terms up to $\phi^{6}$ while the non-supersymmetric potential only has terms up to $\phi^{4}$, the latter still provides a relatively reliable estimate of inflationary quantities.
The main difference between the SUGRA case and the non-supersymmetric case appears when $\phi_{0}$ exceeds $1$. In this region the exponential factor $e^{\phi^{2}/2}$ in the SUGRA case becomes large, which suppresses $r,\,H_{\textrm{inf}}$ and $m_{\phi}$. The energy scales are therefore bounded from above: $H_{\textrm{inf}}<10^{11}\ \textrm{GeV}$ and $m_{\phi}<10^{12}\ \textrm{GeV}$. The non-supersymmetric potential is not able to capture this behavior, which leads to a very different prediction of inflationary observables in this large field scenario.
We further added a SUSY breaking Polonyi sector to the model. If the SUSY breaking scale is much smaller than the Hubble scale, the Polonyi field will stay at the origin and serve as a perturbation to the field. When these two energy scales become comparable, the Polonyi field will move away from the origin and modify the inflation potential. In the opposite extreme the Polonyi field stays close to its post-inflationary value throughout. These effects lead to a nontrivial bound between the SUSY breaking scale and the inflation scale. We find that for a TeV scale SUSY breaking, we need the Hubble scale to be larger than $1$ GeV.
[…]